OpenStudy (anonymous):

Evaluate: 3^2 + (6 − 2) • 4 − 6/3 20 23 50 26

OpenStudy (anonymous):

don't give the answer!!!!

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 9 + 4 * 4 }\) \(\Large \color{purple}{\rightarrow 9 + 16 }\) \(\Large \color{purple}{\rightarrow 25 }\) BODMAS is used.

OpenStudy (diyadiya):

Use PEMDAS

OpenStudy (anonymous):

\[3^2 + 4.4-2 = 9+16-2 =23\]

OpenStudy (diyadiya):

@ParthKohli you've missed something

OpenStudy (anonymous):

ha parth!!!!

OpenStudy (anonymous):